Find the future value of the following ordinary annuity.

Periodic Payment

Payment Interval

Term

Interest Rate

Conversion Period

$1775

6 months

1313

years

9 %

quarterly
Answer #1

Periodic Payment = $ 1773, Payment Interval = 6 months, Tenure = 13 years or (13 x 2) = 26 half-years, Interest Rate = 9%, Compounding Frequency (Conversion Period) = Quarterly,

Applicable Quarterly Rate = 9/4 = 2.25 %

Equivalent Semi-Annual Rate = (1.0225)^(2) - 1 = 0.04551 or 4.551 %

Total Future Value of Periodic Payments = 1773 x (1.04551)^(25) + 1773 x (1.04551)^(24) +..........+ 1773 = 1773 x [{(1.04551)^(26) - 1}/{(1.04551)-1}] = $ 84955.10886 ~ $ 84955.11